OpenStudy (anonymous):

A 0.100-kilogram apple falls from a height of 1.50 meters to 1.00 meters. Ignoring frictional effects, what is the kinetic energy of the apple at this height? 0.00 J 0.49 J 0.98 J 1.47 J

OpenStudy (anonymous):

at the height of 1.5 m, the apple only has potential energy and no kinetic energy since v0=0 m/s so PE=mgh=0.1*9.8*1.5 = 1.47 J , and lets calculate PE at 1 m which is 0.1*9.8*1 = 0.98 J so the difference between them should be KE, which is 0.49 J